OpenStudy (anonymous):

Solve for x: 5^(5x+1) = 25^(x+2)

OpenStudy (anonymous):

Make them both powers of 5 25=5^2 \[5^{5x+1}=(5^{2})^{x+2}\] and then equate the exponents 5x+1=2(x+2) should be easy now